CashMoneyDay:

Identify the axis of symmetry and vertex for the quadratic function x2−2x−15. Axis of Symmetry: x= Vertex: ( , )

CashMoneyDay:

jimthompson5910:

I'll go over a similar example Let's say we had the function y = x^2 - 8x + 12 It's of the form y = ax^2 + bx + c a = 1 b = -8 c = 12 The vertex is of the form (h,k) where h = -b/(2a) h = -(-8)/(2*1) h = 4 This is then plugged into the function to get the y coordinate of the vertex y = x^2 - 8x + 12 y = 4^2 - 8(4) + 12 y = -4 So the vertex is at (h,k) = (4,-4) The axis of symmetry is x = 4 The axis of symmetry is always passing through the vertex, so the x coordinate of the vertex is the same as the axis of symmetry. Keep in mind that this is an example, and the values I got earlier aren't the answers to your specific question. Follow these steps to get the answers you need.